# Compact Cauchy horizons admit constant surface gravity

### v1

@@ 4 Hodge theory and the isometric case
    We recall that if $g_{B}=g_{T}+\chi\otimes\chi$, is the bundle-like metric (i.e. $L_{X}g_{B}=0$ for one and every vector field tangent to the foliation and $\chi$ is positive on the foliation, see e.g. [23, Prop. 2]), then $n$ is defined by the normalization $\chi(n)=1$ and the mean curvature 1-form is $\mu:=i_{n}{\rm d}\chi$. It is basic if $i_{n}\mu=0$, which is always the case, and $i_{n}{\rm d}\mu=0$ which does not necessarily hold. Álvarez López has shown that the basic component $\mu_{b}$ (using orthogonality in the sense of forms induced by $g_{B}$) determines a class (the Álvarez class) $[\mu_{b}]$ which does not depend on the bundle-like metric [1]. In any case, by Dominguez’s theorem [12, 13][56, p. 81], it is possible to choose $\chi$ (and hence $g_{B}$) while leaving $g_{T}$ unaltered, in such a way that the mean curvature 1-form $\mu$ is basic and closed. We define $n$ to be that positive vector such that $\chi(n)=1$ (equivalently, $g_{B}(n,n)=1$), for Dominguez’s choice of $g_{B}$ and $\chi$. Observe that for us it is important that $g_{T}$ does not change [13, Cor. 4.23], as it is the metric induced on the horizon. (Our proof below works using just Álvarez López’s results [1], it is sufficient to work with $\mu_{b}$ instead of $\mu$.)
    From the dominant energy condition, $d\omega(n,.)=0$ and $L_{n}d\omega=0$ so $d\omega$ is a basic 2-form which is closed so, it defines an element in $H_{b}^{2}(\mathcal{F})$. By [9, Prop. III.B.2], we have for the hyperbolic torus $T^{3}_{A}$ endowed with Carrière non-isometric flow, $H_{b}^{2}(\mathcal{F})=0$. So there exists a basic 1-form $\alpha$ such that $d\omega=d\alpha$. This means that $\omega-\alpha$ is a closed 1-form. But the hyperbolic torus has Betti number equal to $1$, this is a well known result, see [18] (paragraph before Theorem 1.1 and the last paragraph of the proof of Proposition 7.2), hence $\textrm{dim}H^{1}(T^{3}_{A})=1$. Since the flow is not isometric, $[\mu](=[\mu_{b}])$ defines a non zero class in $H^{1}_{b}(H)$, see [1, Thm. 6.4][23, Sec. 2.2], and hence in $H^{1}(H)$ (Remark 1.3). Since the dimension of $H^{1}(H)$ is $1$, there exists a constant $\lambda$ and a function $f$ such that $\omega-\alpha=\lambda\mu+df$. Since $\mu$ and $\alpha$ are basic 1-forms, we have $\alpha(n)=\mu(n)=0.$ So $\kappa=\omega(n)=n(f).$ Setting $n^{\prime}=e^{-f}n$ we get $\kappa^{\prime}=0$.
@@ 5 Carrière’s neighborhoods and generic dimension
According to Proposition 4.5 we can focus on compact connected orientable manifolds endowed with a Riemannian flow. Carrière [9] for this type of structure proved that the closures of orbits of a Riemannian flow $\phi$ on a compact connected orientable manifold are tori, and that restricted to each torus, the flow $\phi$ is conjugate (without parameter) to a linear flow. He then established the structure of $\phi$ in a neighborhood of the closure of an orbit [9, Prop. 4].
@@
Proposition 5.1 (Existence of Carrière’s neighborhoods).
@@
where we used the fact that the functions of the partition of unity $\varphi_{i}$ are constant over the orbits. ∎
@@
Proof.
Indeed, let us choose a finite covering $\{V_{i}\}$. From $n(f_{i})=\kappa$ we have multiplying by $\varphi_{i}$, using $n(\varphi_{i})=0$ and summing, $n(f)=\kappa$, where $f=\sum_{i}\varphi_{i}f_{i}$. ∎
Theorem 5.5.
@@
Proof.
By Prop. 5.4 for $\kappa=\omega(n)$, it is sufficient to solve the same problem on a Carrière’s neighborhood $V$. We shall omit the diffeomorphism from $V$ to $[0,1]\times T^{k}\times D^{n-k-1}/\sim$, understanding that $\omega$ and $n$ can be transferred with the diffeomorphism if needed. At no step of the argument we shall need the metric $g_{T}$ (which might be non-unique). We denote with $n$ the original vector field and with $n^{\prime}=\frac{\partial}{\partial s}$ the natural vector field of the suspension where $s$ is the parameter of $[0,1]$. This vector field is defined just on $V$.
However, its restriction to $V_{1/2}:=[0,1]\times T^{k}\times D_{1/2}^{n-k-1}/\sim$, where $D_{1/2}^{n-k-1}$ is the open ball of radius 1/2 of $\mathbb{R}^{n-k-1}$, can be easily extended to the whole $H$ (e.g. by taking a convex combination with another global choice, making use of function $\rho$, c.f. Remark 5.2). We denote this extension with $n^{\prime\prime}$. We recall that the properties of Theorem 3.1 do not depend on the vector field (Remark 3.2), thus they hold also using $n^{\prime\prime}$ which implies $\lim_{s\to\infty}\frac{1}{s}\int_{0}^{s}\omega(\dot{x}(s)){\rm d}s=0$, where $x$ is any orbit on $V_{1/2}$ parametrized with the natural suspension parameter $s$.
Let $L$ be the closure of orbit corresponding to $r=0$. Let $\rho$ be the function on the ball $D^{n-k-1}$ constructed in Remark 5.2. We already observed that the region $r<1/2$, $r:=\|y\|$, provides another Carrière’s neighborhood of $L$ denoted $V_{1/2}$ as it is the suspension with base $T^{k}\times D_{1/2}^{n-k-1}$. The 1-form $\omega^{\prime}:=\rho\omega$ vanishes for $r>1-\epsilon$ and is coincident with $\omega$ for $r<1/2$ hence in $V_{1/2}$. Moreover, since $\rho$ is constant over every orbit in $V$, $i_{n}{\rm d}\omega^{\prime}=(i_{n}{\rm d}\rho)\omega+\rho i_{n}{\rm d}\omega=0$.
@@
Formally, the map does not change, and it is still an isometry because the rotation preserves the function $\Omega$. This isometric property implies that it induces a Riemannian flow with tangent field $\partial/\partial s$ which we keep denoting $n^{\prime}$. In conclusion, the Carrière neighborhood can be regarded as a subset of a compact connected manifold $N$ which is obtained via a suspension. Coming from a suspension, the flow is isometric, in fact the direct sum of metrics $h={\rm d}t^{2}\oplus g_{\mathbb{T}^{k}}\oplus g_{S^{n-k-1}_{\textrm{def}}}$ is such that $L_{n}h=0$ (see the discussion in [23]). Furthermore, the orientability of $T^{k}\times S^{n-k-1}_{\textrm{def}}$ and the fact that $O$ has unit determinant implies that $N$ is orientable. Now, we extend $\omega$ further, defining $\omega^{\prime}=0$ in $N\backslash V$. Then, $N$, $n^{\prime}$, $h$, $\omega^{\prime}$, satisfy the assumptions of Theorem 4.4 thus there are a smooth function $f:N\to\mathbb{R}$ and a constant $c$ such that $n^{\prime}(f)=\omega^{\prime}(n^{\prime})-c$. We denote with $\check{f}$ the restriction to $V_{1/2}$. As there $\omega^{\prime}=\omega$ we get $n^{\prime}(\check{f})=\omega(n^{\prime})-c$. Integrating this equation and using $\lim_{s\to\infty}\frac{1}{s}\int_{0}^{s}\omega(\dot{x}(s)){\rm d}s=0$ and $n^{\prime}=\frac{{\rm d}}{{\rm d}s}$ we get $c=0$ (because $f$, and hence $\check{f}$, is bounded as it is a continuous function over the compact set $N$). Using the linearity with respect to $n$ we arrive at $n(\check{f})=\omega(n)$ on $V_{1/2}$. As the equation can be solved over Carrière neighborhoods that cover the whole $H$, by passing to a finite covering we get a global solution on $H$. ∎
@@
Proof of Theorem 1.6.
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the dominant energy condition. Let $H$ be a compact degenerate horizon, we want to prove that there exists a smooth future-directed lightlike tangent vector field $n$ of zero surface gravity, $\omega(n)=0$, where $\nabla_{X}n=\omega(X)n$.
By the dominant energy condition $i_{n}{\rm d}\omega=0$ (Prop. 2.2). By Prop. 4.5 we can assume, without loss of generality, that $H$ is orientable. By Theorem 3.3 the equivalent properties of Theorem 3.1 are satisfied. By Theorem 5.5 there is a smooth function such that $n(f)=\omega(n)$, thus by Eq. (3) $n^{\prime}=e^{-f}n$ has zero surface gravity. ∎
@@ 6 Further results in the torus case
Theorem 6.3.
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the dominant energy condition. Let $H$ be a compact degenerate horizon, and suppose that it is diffeomorphic to the torus $T^{s}$, the Riemannian flow being conjugate (without parameter) to a linear flow on the torus with dense orbits. Then there exists a smooth future-directed lightlike tangent vector field $n$ of zero surface gravity, namely geodesic, and moreover with $\omega$ defined as in Eq. (1), we have ${\rm d}\omega=0$. That is, $\omega\in H^{1}_{b}(H)$.
Proof.
We already know that we can find $n$ of zero surface gravity, so let $\omega$ be the associated 1-form from Eq. (1) so that $\omega(n)=0$. By the dominant energy condition $i_{n}{\rm d}\omega=0$ thus $\omega$ is a basic 1-form. Note that $n$ is proportional to the vector field $n^{\prime}$ mentioned in the statement of Theorem 6.1 (and there denoted $n$) thus $i_{n^{\prime}}{\rm d}\omega=0$, thus from that theorem we get ${\rm d}\omega=0$. ∎
@@ 7 Conclusions
We proved that, in any spacetime dimension and under the dominant energy condition, every smooth connected totally geodesic compact null hypersurface (including compact Cauchy horizons) admits a smooth, future-directed, lightlike vector field $n$ of constant surface gravity. That is, we resolved the open degenerate case by establishing the existence of a geodesic lightlike field whenever a complete generator exists. This showed that in the degenerate case, for a suitable choice of $n$, the 1-form $\omega$ is basic. The proof uses techniques from ergodic theory, Hodge theory and the theory of Riemannian flows.

### v2

@@ 4 Hodge theory and the isometric case
    We recall that if $g_{B}=g_{T}+\chi\otimes\chi$, is the bundle-like metric (i.e. $L_{X}g_{B}=0$ for one and every vector field tangent to the foliation and $\chi$ is positive on the foliation, see e.g. (minguzzi24c, , Prop. 2)), then $n$ is defined by the normalization $\chi(n)=1$ and the mean curvature 1-form is $\mu:=i_{n}{\rm d}\chi$. It is basic if $i_{n}\mu=0$, which is always the case, and $i_{n}{\rm d}\mu=0$ which does not necessarily hold. Álvarez López has shown that the basic component $\mu_{b}$ (using orthogonality in the sense of forms induced by $g_{B}$) determines a class (the Álvarez class) $[\mu_{b}]$ which does not depend on the bundle-like metric alvarez92 . In any case, by Dominguez’s theorem dominguez95 ; dominguez98 (tondeur97, , p. 81), it is possible to choose $\chi$ (and hence $g_{B}$) while leaving $g_{T}$ unaltered, in such a way that the mean curvature 1-form $\mu$ is basic and closed. We define $n$ to be that positive vector such that $\chi(n)=1$ (equivalently, $g_{B}(n,n)=1$), for Dominguez’s choice of $g_{B}$ and $\chi$. Observe that for us it is important that $g_{T}$ does not change (dominguez98, , Cor. 4.23), as it is the metric induced on the horizon. (Our proof below works using just Álvarez López’s results alvarez92 , it is sufficient to work with $\mu_{b}$ instead of $\mu$.)
    From the null energy condition, $d\omega(n,.)=0$ and $L_{n}d\omega=0$ so $d\omega$ is a basic 2-form which is closed so, it defines an element in $H_{b}^{2}(\mathcal{F})$. By (carriere84, , Prop. III.B.2), we have for the hyperbolic torus $T^{3}_{A}$ endowed with Carrière non-isometric flow, $H_{b}^{2}(\mathcal{F})=0$. So there exists a basic 1-form $\alpha$ such that $d\omega=d\alpha$. This means that $\omega-\alpha$ is a closed 1-form. But the hyperbolic torus has Betti number equal to $1$, this is a well known result, see funar13 (paragraph before Theorem 1.1 and the last paragraph of the proof of Proposition 7.2), hence $\textrm{dim}H^{1}(T^{3}_{A})=1$. Since the flow is not isometric, $[\mu](=[\mu_{b}])$ defines a non zero class in $H^{1}_{b}(H)$, see (alvarez92, , Thm. 6.4)(minguzzi24c, , Sec. 2.2), and hence in $H^{1}(H)$ (Remark 1). Since the dimension of $H^{1}(H)$ is $1$, there exists a constant $\lambda$ and a function $f$ such that $\omega-\alpha=\lambda\mu+df$. Since $\mu$ and $\alpha$ are basic 1-forms, we have $\alpha(n)=\mu(n)=0.$ So $\kappa=\omega(n)=n(f).$ Setting $n^{\prime}=e^{-f}n$ we get $\kappa^{\prime}=0$.
@@ 5 Carrière’s neighborhoods and generic dimension
According to Proposition 3 we can focus on compact connected orientable manifolds endowed with a Riemannian flow. Carrière carriere84 for this type of structure proved that the closures of orbits of a Riemannian flow $\phi$ on a compact connected orientable manifold are tori, and that restricted to each torus, the flow $\phi$ is conjugate (without parameter) to a linear flow. He then established the structure of $\phi$ in a neighborhood of the closure of an orbit (carriere84, , Prop. 4).
@@
Proposition 4 (Existence of Carrière’s neighborhoods)
@@
where we used the fact that the functions of the partition of unity $\varphi_{i}$ are constant over the orbits.
@@
Proof
Indeed, let us choose a finite covering $\{V_{i}\}$. From $n(f_{i})=\kappa$ we have multiplying by $\varphi_{i}$, using $n(\varphi_{i})=0$ and summing, $n(f)=\kappa$, where $f=\sum_{i}\varphi_{i}f_{i}$.
Theorem 5.1
@@
Proof
By Prop. 6 for $\kappa=\omega(n)$, it is sufficient to solve the same problem on a Carrière’s neighborhood $V$. We shall omit the diffeomorphism from $V$ to $[0,1]\times T^{k}\times D^{n-k-1}/\sim$, understanding that $\omega$ and $n$ can be transferred with the diffeomorphism if needed. At no step of the argument we shall need the metric $g_{T}$ (which might be non-unique). We denote with $n$ the original vector field and with $n^{\prime}=\frac{\partial}{\partial s}$ the natural vector field of the suspension where $s$ is the parameter of $[0,1]$. This vector field is defined just on $V$.
However, its restriction to $V_{1/2}:=[0,1]\times T^{k}\times D_{1/2}^{n-k-1}/\sim$, where $D_{1/2}^{n-k-1}$ is the open ball of radius 1/2 of $\mathbb{R}^{n-k-1}$, can be easily extended to the whole $H$ (e.g. by taking a convex combination with another global choice, making use of function $\rho$, c.f. Remark 4). We denote this extension with $n^{\prime\prime}$. We recall that the properties of Theorem 3.1 do not depend on the vector field (Remark 3), thus they hold also using $n^{\prime\prime}$ which implies $\lim_{s\to\infty}\frac{1}{s}\int_{0}^{s}\omega(\dot{x}(s)){\rm d}s=0$, where $x$ is any orbit on $V_{1/2}$ parametrized with the natural suspension parameter $s$.
Let $L$ be the closure of orbit corresponding to $r=0$. Let $\rho$ be the function on the ball $D^{n-k-1}$ constructed in Remark 4. We already observed that the region $r<1/2$, $r:=\|y\|$, provides another Carrière’s neighborhood of $L$ denoted $V_{1/2}$ as it is the suspension with base $T^{k}\times D_{1/2}^{n-k-1}$. The 1-form $\omega^{\prime}:=\rho\omega$ vanishes for $r>1-\epsilon$ and is coincident with $\omega$ for $r<1/2$ hence in $V_{1/2}$. Moreover, since $\rho$ is constant over every orbit in $V$, $i_{n}{\rm d}\omega^{\prime}=(i_{n}{\rm d}\rho)\omega+\rho i_{n}{\rm d}\omega=0$.
@@
Formally, the map does not change, and it is still an isometry because the rotation preserves the function $\Omega$. This isometric property implies that it induces a Riemannian flow with tangent field $\partial/\partial s$ which we keep denoting $n^{\prime}$. In conclusion, the Carrière neighborhood can be regarded as a subset of a compact connected manifold $N$ which is obtained via a suspension. Coming from a suspension, the flow is isometric, in fact the direct sum of metrics $h={\rm d}t^{2}\oplus g_{\mathbb{T}^{k}}\oplus g_{S^{n-k-1}_{\textrm{def}}}$ is such that $L_{n}h=0$ (see the discussion in minguzzi24c ). Furthermore, the orientability of $T^{k}\times S^{n-k-1}_{\textrm{def}}$ and the fact that $O$ has unit determinant implies that $N$ is orientable. Now, we extend $\omega$ further, defining $\omega^{\prime}=0$ in $N\backslash V$. Then, $N$, $n^{\prime}$, $h$, $\omega^{\prime}$, satisfy the assumptions of Theorem 4.3 thus there are a smooth function $f:N\to\mathbb{R}$ and a constant $c$ such that $n^{\prime}(f)=\omega^{\prime}(n^{\prime})-c$. We denote with $\check{f}$ the restriction to $V_{1/2}$. As there $\omega^{\prime}=\omega$ we get $n^{\prime}(\check{f})=\omega(n^{\prime})-c$. Integrating this equation and using $\lim_{s\to\infty}\frac{1}{s}\int_{0}^{s}\omega(\dot{x}(s)){\rm d}s=0$ and $n^{\prime}=\frac{{\rm d}}{{\rm d}s}$ we get $c=0$ (because $f$, and hence $\check{f}$, is bounded as it is a continuous function over the compact set $N$). Using the linearity with respect to $n$ we arrive at $n(\check{f})=\omega(n)$ on $V_{1/2}$. As the equation can be solved over Carrière neighborhoods that cover the whole $H$, by passing to a finite covering we get a global solution on $H$.
@@
Proof (Proof of Theorem 1.4)
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the null energy condition. Let $H$ be a compact degenerate horizon, we want to prove that there exists a smooth future-directed lightlike tangent vector field $n$ of zero surface gravity, $\omega(n)=0$, where $\nabla_{X}n=\omega(X)n$.
By the null energy condition $i_{n}{\rm d}\omega=0$ (Cor. 1). By Prop. 3 we can assume, without loss of generality, that $H$ is orientable. By Theorem 3.2 the equivalent properties of Theorem 3.1 are satisfied. By Theorem 5.1 there is a smooth function such that $n(f)=\omega(n)$, thus by Eq. (3) $n^{\prime}=e^{-f}n$ has zero surface gravity.
@@ 6 Further results in the torus case
Theorem 6.2
Let $(M,g)$ be a smooth spacetime (of any dimension) and assume the null energy condition. Let $H$ be a compact degenerate horizon, and suppose that it is diffeomorphic to the torus $T^{s}$, the Riemannian flow being conjugate (without parameter) to a linear flow on the torus with dense orbits. Then there exists a smooth future-directed lightlike tangent vector field $n$ of zero surface gravity, namely geodesic, and moreover with $\omega$ defined as in Eq. (1), we have ${\rm d}\omega=0$. That is, $\omega\in H^{1}_{b}(H)$.
Proof
We already know that we can find $n$ of zero surface gravity, so let $\omega$ be the associated 1-form from Eq. (1) so that $\omega(n)=0$. By the null energy condition $i_{n}{\rm d}\omega=0$ thus $\omega$ is a basic 1-form. Note that $n$ is proportional to the vector field $n^{\prime}$ mentioned in the statement of Theorem 6.1 (and there denoted $n$) thus $i_{n^{\prime}}{\rm d}\omega=0$, thus from that theorem we get ${\rm d}\omega=0$.
@@ 7 Conclusions
We proved that, in any spacetime dimension and under the null energy condition, every smooth connected totally geodesic compact null hypersurface (including compact Cauchy horizons) admits a smooth, future-directed, lightlike vector field $n$ of constant surface gravity. That is, we resolved the open degenerate case by establishing the existence of a geodesic lightlike field whenever a complete generator exists. This showed that in the degenerate case, for a suitable choice of $n$, the 1-form $\omega$ is basic. The proof uses techniques from ergodic theory, Hodge theory and the theory of Riemannian flows.
@@ Acknowledgments
Declarations
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflicts of Interest
The authors have no relevant financial or non-financial interests to disclose.